# Localized Band Gap Resonances in Defect-embedded Monoatomic Chains

### v1

@@ II.3 Multi-defect embedded grounded monoatomic chain
Now, consider the configurations described by the y-axis of the contour plots in Fig. 4a that represents a Multi-DMoA with an edge ($x_{\mathrm{d1}}=0$) and a bulk ($x_{\mathrm{d2}}\in[1,\mathrm{N}-2]$) defect. Fig. 4b plots the numerical edge and bulk defect frequencies, $\Omega_{\mathrm{d1}}$ and $\Omega_{\mathrm{d2}}$, respectively, for different locations of the bulk defect, $x_{\mathrm{d2}}$ in the Multi-DMoA. Consistent with the error plots, both $\Omega_{\mathrm{d1}}$ and $\Omega_{\mathrm{d2}}$ asymptote to the respective edge-bulk analytical estimates as the bulk defect migrates farther from the edge defect. To further explore interesting dynamics, we consider the Multi-DMoA described by the $\times$ -marker in Fig. 4a,b. Fig. 4c-i (Fig. 4c-ii) plots the FRF when the edge (bulk) defect is subject to a harmonic excitation and its displacement monitored. The relatively low amplitude and monotonic FRF curves when exciting the edge (bulk) defect mass at the other defect frequency, $\Omega_{\mathrm{d2}}$ ($\Omega_{\mathrm{d1}}$) presents an interesting insight in Fig. 4c-i (Fig. 4c-ii). The edge and bulk defect mode shapes plotted in Fig. 4c-iii explain this peculiar dynamical behavior of the Multi-DMoA. In the bulk defect mode, $x_{\mathrm{d2}}=13$ constitutes an anti-node, while $x_{\mathrm{d1}}=0$ constitutes a node. Consequently, exciting the edge defect mass ($x_{\mathrm{d1}}=0$) at $\Omega_{\mathrm{d2}}$ fails to excite the bulk defect mode, and exciting the bulk defect mass ($x_{\mathrm{d2}}=13$) at $\Omega_{\mathrm{d1}}$ fails to excite the edge defect mode, leading to a relatively muted dynamical response at the respective excitation locations in both scenarios. Therefore, the analytical approach provides a robust estimate of the defect mode frequencies, mode shapes, and the velocity amplitude envelope even in the presence of multiple defects in the DMoA, provided the additional defects are nodal points in the eigenmode associated with the considered defect.
III Defect-embedded Acoustic Metamaterials
@@
In this section, we demonstrate how tailored edge DMoA-based AMMs can – (i) overcome the absence of an inertial frame of reference by accommodating an in-house virtual ground within the AMM structure, and (ii) manifest a customized amplitude and narrow-band frequency in response to broadband acoustic fields, e.g., acoustic waves generated due to vehicular traffic, dynamic pressure field of a fluid which may be relevant to applications, e.g., noise/vibration mitigation, energy harvesting, flow control.
@@
Note, we include nominal viscous damping, $c_{\mathrm{r}}=c/\sqrt{km}=0.01$ in numerical simulations presented in this section, to realize bounded acoustic responses and test the robustness of our analytical model.
III.1 Acoustic Metamaterial with a virtual ground
@@
Figure 5: Acoustic Metamaterial (AMM) with a virtual ground. (a-i) Schematics of AMMs featuring a rigid inertial and a non-inertial ground. (a-ii) FFTs of velocity response, $\dot{v}_{\mathrm{d}}(\bar{t})$ in an AMM with rigid inertial grounding (red) and non-inertial grounding (blue). Both AMMs are subject to force $\mathrm{F}(\bar{t})$ (green), however, the rigid grounding case has no environmental noise while the non-inertial grounding case features a broadband environmental noise (orange) contaminating the AMM response via the grounding springs. (b-i) Schematic of a modified AMM where the last mass functions as a virtual ground, and isolates the AMM from undesirable external noise while simultaneously retaining the edge defect mode characteristics relatively intact. (b-ii) FFTs of $\dot{v}_{\mathrm{d}}(\bar{t})$ in an AMM with virtual grounding subject to force $\mathrm{F}(\bar{t})$. (c-i) Reduced order mass-spring model of the AMM with virtual grounding, where grounding springs connect all masses to the last mass (opposite boundary), (c-ii) eigenmode of the AMM at $\Omega_{\mathrm{d,AMM}}=0.8447$, and (c-iii) the convergence plot of the edge defect frequency, $\Omega_{\mathrm{d,AMM}}$.
@@
As described in Sec. II, although grounding elements are not necessary for the existence of defect modes, their presence enables the existence of a defect mode as the principal eigenmode of a DMoA-based AMM within the bounded BG – $\Omega\in[0,\sqrt{k_{\mathrm{g,r}}})$. However, AMMs have been deployed in a variety of environments, e.g., rotating shafts[47], aerial vehicles[41], seismic structures[48] where a true inertial frame of reference may not be accessible. In this context, we present a modified edge DMoA-based AMM that can function unhindered without a rigid inertial reference by utilizing one of its boundaries as a virtual ground.
@@
The schematics in Fig. 5(a,b-i) contrast AMM configurations with an inertial, non-inertial, and virtual ground. Fig. 5(a,b-ii) contrasts the FFTs of the edge defect velocities when subject to a Gaussian wave-packet type forcing, $\mathrm{F}(\bar{t})$ (centered at $\Omega_{\mathrm{d,DMoA}}$) in these three grounding configurations. The FFT (red) in the rigid grounding scenario reveals a behavior consistent with dynamics characterized by the edge DMoA model in Fig. 1b. However, the presence of broadband environmental noise (orange) renders the grounding mechanism non-inertial, causing the FFT (blue) in this scenario to deviate from its rigid grounding counterpart. Even though the environmental noise is chosen to have a maximum amplitude an order lower ($O(10^{-1})$) than the excitation signal, $\mathrm{F}(\bar{t})$, the impact on the AMM dynamics is evident.
@@
Since the nature of the external grounding reference is dictated by the operating environment, we consider an alternate AMM configuration as shown in Fig. 5(b-i) to mitigate the impact of a non-inertial grounding reference. In the modified AMM, all constituent masses are now connected to the last mass which functions as an in-house virtual ground, isolating all AMM masses from the external environmental noise. Fig. 5(c-i) presents an equivalent mass-spring reduced order model for this altered AMM configuration with a virtual ground. An eigen-analysis (see Sec. A5) reveals completely new eigenvalues compared to the edge DMoA-based AMM, with three BG resonances – $\Omega=0$ and $\Omega_{\mathrm{d,AMM}}=0.8447$ in the low-frequency BG, and $\Omega=6.4095$ in the high-frequency BG, and the remainder lying within the PB range of the base MoA. Resonances at $\Omega=0$ and $\Omega_{\mathrm{d,AMM}}=0.8447$, respectively, correspond to the rigid body translation mode and the localized mode shape plotted in Fig. 5(c-ii) while the resonance at $\Omega=6.4095$ corresponds to a new eigenmode localized at the last mass, created as a result of the altered grounding spring connections. Though the eigenmode frequencies and shapes are significantly different from those of an edge DMoA-based AMM, the eigenmode at $\Omega_{\mathrm{d,AMM}}=0.8447$ (Fig. 5(c-ii)) emerges as an exception, existing relatively unaltered compared to the edge defect mode shown in the inset of Fig. 1b, owing to the last mass being a nodal point in the defect mode. The FFT curve (red) in Fig. 5(b-ii) captures the AMM dynamics in the presence of an in-house virtual ground which is identical to that of an AMM with a rigid ground (Fig. 5(a-ii)). Therefore, our modified analytical approach in conjunction with the presence of an in-house virtual ground provides an enhanced capability to design a DMoA-based AMM that can function efficiently at a prescribed BG resonance regardless of the nature of the operating environment.
@@
A closer look at Fig. 5(b-ii) reveals a slight change in defect modal frequency, $\Omega_{\mathrm{d,AMM}}$ compared to the analytical DMoA edge defect frequency, $\Omega_{\mathrm{d,DMoA}}=0.8165$. This discrepancy stems from the limitation that the decaying eigenmode reaches a near-zero and not an absolute zero displacement at the last mass, i.e., $v_{\mathrm{N}}\approx 0$. Fig. 5(c-iii) illustrates this phenomenon, where choosing a larger AMM (large $\mathrm{N}$) aids in converging the displacement, $v_{\mathrm{N}}\rightarrow 0$ and the frequency, $\Omega_{\mathrm{d,AMM}}\rightarrow\Omega_{\mathrm{d,DMoA}}$.
III.2 Customized Acoustic absorber

### v2

@@ 2.4 Multi-defect embedded grounded monoatomic phononic crystal
Now, consider the configurations described by the y-axis of the contour plots in Fig. 4a that represents a Multi-DMoA PnC with an edge ($x_{\mathrm{d1}}=0$) and a bulk ($x_{\mathrm{d2}}\in[1,\mathrm{N}-2]$) defect. Fig. 4b plots the numerical edge and bulk defect frequencies, $\Omega_{\mathrm{d1}}$ and $\Omega_{\mathrm{d2}}$, respectively, for different locations of the bulk defect, $x_{\mathrm{d2}}$ in the Multi-DMoA PnC. Consistent with the error plots, both $\Omega_{\mathrm{d1}}$ and $\Omega_{\mathrm{d2}}$ asymptote to the respective edge-bulk analytical estimates as the bulk defect migrates farther from the edge defect. To further explore interesting dynamics, we consider the Multi-DMoA PnC described by the $\times$ -marker in Fig. 4a,b. Fig. 4c-i (Fig. 4c-ii) plots the FRF when the edge (bulk) defect is subject to a harmonic excitation and its displacement monitored. The relatively low amplitude and monotonic FRF curves when exciting the edge (bulk) defect mass at the other defect frequency, $\Omega_{\mathrm{d2}}$ ($\Omega_{\mathrm{d1}}$) presents an interesting insight in Fig. 4c-i (Fig. 4c-ii). The edge and bulk defect mode shapes plotted in Fig. 4c-iii explain this peculiar dynamical behavior of the Multi-DMoA PnC. In the bulk defect mode, $x_{\mathrm{d2}}=13$ constitutes an anti-node, while $x_{\mathrm{d1}}=0$ constitutes a node. Consequently, exciting the edge defect mass ($x_{\mathrm{d1}}=0$) at $\Omega_{\mathrm{d2}}$ fails to excite the bulk defect mode, and exciting the bulk defect mass ($x_{\mathrm{d2}}=13$) at $\Omega_{\mathrm{d1}}$ fails to excite the edge defect mode, leading to a relatively muted dynamical response at the respective excitation locations in both scenarios. Therefore, the analytical approach provides a robust estimate of the defect mode frequencies, mode shapes, and the velocity amplitude envelope even in the presence of multiple defects in the DMoA PnC, provided the additional defects are nodal points in the eigenmode associated with the considered defect.
2.5 Defect modes in 2D DMoA PnCs
This section proposes a solution approach to quantify the defect resonance frequency and mode shape in a 2D DMoA PnC. Consider the non-dimensional equations of motion of a uniform 2D membrane (Fig. 5(a-i)) modeled as a 2D MoA PnC:
$v_{,\bar{t}\bar{t}}^{(p,q)}+\left(4v^{(p,q)}-v^{(p+1,q)}-v^{(p-1,q)}-v^{(p,q+1)}-v^{(p,q-1)}\right)+k_{\mathrm{g,r}}v^{(p,q)}=0,$ (5)
where, $v^{(p,q)}$ represents the out-of-plane displacement of the mass located in the $p-$th column and $q-$th row. The dispersion relations for the 2D membrane, assuming a plane wave ansatz, $v^{(p,q)}=\tilde{v}^{(p,q)}e^{i(\mathbf{k}\cdot\mathbf{x}-\Omega\bar{t})}\ \forall\ \mathbf{k}=[\kappa_{p},\kappa_{q}]^{\mathrm{T}},\mathrm{x}=[x_{p},x_{q}]^{\mathrm{T}}$, where $x_{j}=j-1$ for $j=p,q$, can be calculated as:
$\Omega(\mathbf{k})=\sqrt{k_{\mathrm{g,r}}+4\sin^{2}\left({\frac{\kappa_{p}}{2}}\right)+4\sin^{2}\left({\frac{\kappa_{q}}{2}}\right)}$ (6)
where, $\mathrm{Re}\left[\kappa_{p,q}\right]\in[0,\pi]$. Fig. 5(b-i) plots the dispersion curve calculated in Eq. (6) using the primitive unit cell (Fig. 5(a-i)), associated with the irreducible Brillouin zone.
Defect modes in 2D DMoA PnCs can be explored in three distinct configurations – (i) a corner mode, (ii) an edge mode, and (iii) a bulk mode[32]. Though the direct application of the proposed analysis method (Eq. (3)) to 2D lattices can become cumbersome, we can develop a new method to analyze the defect modes in 2D DMoA PnCs taking insights from the 1D analysis.
Consider the equations of motion of a defect mass in the 2D DMoA PnC:
$m_{\mathrm{r}}v_{,\bar{t}\bar{t}}^{(p,q)}+k_{\mathrm{r}}\left[n_{c}v^{(p,q)}-H\left(v^{(p^{\prime},q^{\prime})}\right)\right]+k_{\mathrm{gd,r}}v^{(p,q)}=0,$ (7)
where, $n_{c}$, and the function, $H$ represent the coordination number and the connectivity function, respectively, for different defect configurations listed in Tab. 1.
Types of Defect
Corner Edge Bulk
Left bottom, Right bottom, Right top, Left top, Bottom, Right, Top, Left,
$p=1$ $p=\mathrm{N}$, $p=\mathrm{N}$ $p=1$, $p\in\mathcal{S}$ $p=\mathrm{N}$ $p\in\mathcal{S}$, $p=1$, $p\in\mathcal{S}$
$q=1$ $q=1$ $q=\mathrm{N}$ $q=\mathrm{N}$ $q=1$ $q\in\mathcal{S}$ $q=\mathrm{N}$ $q\in\mathcal{S}$ $q\in\mathcal{S}$
$n_{c}$ 2 2 2 2 3 3 3 3 4
$H\left(v^{(p^{\prime},q^{\prime})}\right)$ $v^{(p+1,q)}$
$v^{(1,2)}$ $v^{(\mathrm{N}-1,1)}$ $v^{(\mathrm{N}-1,\mathrm{N})}$ $v^{(1,\mathrm{N}-1)}$ $v^{(p+1,1)}$ $v^{(\mathrm{N},q+1)}$ $v^{(p+1,\mathrm{N})}$ $v^{(1,q+1)}$ $+v^{(p-1,q)}$
$+v^{(2,1)}$ $+v^{(\mathrm{N},2)}$ $+v^{(\mathrm{N},\mathrm{N}-1)}$ $+v^{(2,\mathrm{N})}$ $+v^{(p-1,1)}$ $+v^{(\mathrm{N},q-1)}$ $+v^{(p-1,\mathrm{N})}$ $+v^{(1,q-1)}$ $+v^{(p,q+1)}$
$+v^{(p,2)}$ $+v^{(\mathrm{N}-1,q)}$ $+v^{(p,\mathrm{N}-1)}$ $+v^{(2,q)}$ $+v^{(p,q-1)}$
Table 1: Coordination number and interaction spring connections. $\mathcal{S}=[2,\mathrm{N}-1]$
The continuum approximation for Eq. (5) yields:
$v_{,\bar{t}\bar{t}}-\Delta v+k_{\mathrm{g,r}}v=0,$
where, $\Delta=\frac{\partial^{2}}{\partial x_{p}^{2}}+\frac{\partial^{2}}{\partial x_{q}^{2}}$ is the Laplacian in the Cartesian plane, $x_{p}-x_{q}$. The 2D MoA PnC arrangement is such that, a single material defect at any lattice location will manifest an axis-symmetric decay. Assuming axis-symmetric conditions, the Laplacian can be transformed into polar coordinates: $\Delta=\frac{\partial^{2}}{\partial r^{2}}+\frac{1}{r}\cdot\frac{\partial}{\partial r}$. For a lattice location sufficiently far away, i.e., $r\rightarrow\infty$, from the defect location ($r=0$), the above equation becomes:
$v_{,\bar{t}\bar{t}}-\frac{\partial^{2}v}{\partial r^{2}}+k_{\mathrm{g,r}}v=0,$
resembling the continuum approximation for a discrete 1D MoA PnC. Therefore, the axis-symmetric dispersion relation for waves traveling radially in a discrete lattice is given by Eq. (2).
From our 1D analysis in Sec. 2.1, we observe that the defect frequencies and decay rates, $\{\Omega_{\mathrm{d}},\kappa_{I}\}$ of 1D DMoA PnCs satisfy the dispersion relations of the underlying MoA PnC (Eq.(2), Fig. 1(a-i)). Informed by this observation, we combine the axis-symmetric dispersion (Eq. (2)) and the equations of motion of the defect mass (Eq. (7)), to predict the defect mode characteristics in 2D DMoA PnCs (Fig. 5(a-ii)).
Given a radial decay rate, $\kappa_{I}$, the displacement of every mass connected to the defect mass can be given as:
$v^{(p^{\prime},q^{\prime})}=v^{(p,q)}e^{-\kappa_{I}},$ (8)
rendering $H\left(v^{(p^{\prime},q^{\prime})}\right)=n_{c}v^{(p,q)}e^{-\kappa_{I}}$. Substituting the ansatz $v^{(p,q)}=\widetilde{v}_{\mathrm{d}}e^{i\Omega_{\mathrm{d}}\bar{t}}$ in Eq. (8) and combining it with Eq. (2), and Eq. (7) , we get:
$m_{\mathrm{r}}X+\frac{(m_{\mathrm{r}}-k_{\mathrm{r}}n_{c})}{X}+2\delta=0,$ (9)
where,
$X=e^{\kappa_{I}},\delta=\frac{(k_{\mathrm{gd,r}}-m_{\mathrm{r}}k_{\mathrm{g,r}})+(k_{\mathrm{r}}n_{c}-2m_{\mathrm{r}})}{2}$
Rewriting Eq. (9) in a quadratic form and estimating the roots, we get:
$\kappa_{I}=\log\left(\frac{-\delta+\sqrt{\delta^{2}-m_{\mathrm{r}}(m_{\mathrm{r}}-k_{\mathrm{r}}n_{c})}}{m_{\mathrm{r}}}\right),$ (10a)
$\Omega_{\mathrm{d}}=\sqrt{\frac{k_{\mathrm{gd,r}}+k_{\mathrm{r}}n_{c}(1-e^{-\kappa_{I}})}{m_{\mathrm{r}}}},$ (10b)
Eq. (10b) represents the analytical decay rates and the defect mode frequencies in a 2D DMoA PnC lattice with any of the three types of defects.
Consider a 2D Multi-DMoA PnC ($\mathrm{N}_{p}=\mathrm{N}_{q}=21$) shown in Fig. 5(a-ii), embedded with a corner ($x_{p,q}=0$), an edge ($x_{p}=0,x_{q}=11$) and a bulk ($x_{p,q}=11$) defect. Applying Eq. (10b), we calculate the modal characteristics for each of the three defect resonances in the 2D lattice. Fig. 5(b-ii,iii) contrasts the analytical and numerical values of $\{\Omega_{\mathrm{d}},\kappa_{I}\}$ along both the in-plane ($x_{p,q}$) directions, and Fig. 5c plots the mode shapes of the respective BG resonances. The numerical and analytical $\Omega_{\mathrm{d}}$ values show an excellent agreement with an average error of $1.94\%$ between the three modes. The analytical decay rate also correlates well with the in-plane identical numerical decay rates, with a nominally higher average error of $5.23\%$ between the three modes. The analytical solutions have been tested for various material parameters and, show good correlation with their numerical counterparts with a low average error ($\sim 1\%-8\%$). Hence, Eq. (10b) presents a valuable analytical form of the modal characteristics of singular defects in a 2D DMoA PnC, that can be successfully extended to 2D Multi-DMoA PnCs (similar to the 1D Multi-DMoA PnC, Sec. 2.4) provided the embedded defects lie outside the proximal influence zones of each other.
3 Applications of Defect-embedded PnCs
@@
In this section, we demonstrate how tailored edge DMoA-based PnCs can – (i) overcome the absence of an inertial frame of reference by accommodating an in-house virtual ground within the PnC structure, and (ii) manifest a customized amplitude and narrow-band frequency in response to broadband acoustic fields.
Engineering a phononic BG resonance[38, 50] in PnCs has gained relevance in recent years in passive flow control applications, which brings together two rich fields of research - fluid dynamics, and PnCs and metamaterials. This presents an exciting opportunity for the two communities to collaborate and develop novel fluid-structure interaction strategies to achieve beneficial outcomes, e.g., fluid flow control using passive PnC designs, and efficiently harvesting energy from diverse fluid flows. In this context, the virtual ground design is useful in flow control applications where the PnC needs to be deployed in non-inertial environments, e.g., wings of moving aerial vehicles[41], underwater vehicles. An in-house virtual ground keeps the PnC isolated from external noise, e.g., wing vibrations, and engine noise, allowing effective airflow-PnC interaction leading to a significant reduction in operational and fuel costs by delaying laminar-to-turbulent transition[38], reducing turbulent drag[51]. Additionally, the relatively simple dynamical equations of DMoA-based PnCs and the existence of analytical relations between the defect mode characteristics and material properties, compared to other BG resonances (e.g., truncation resonances, topological modes) presents a favorable strategy to rapidly generate, test, and iterate PnC designs for flow control.
The customized acoustic absorber design is relevant for noise/vibration mitigation, and energy harvesting applications in environments subsumed in broadband acoustic noise, e.g., vehicular traffic, and machine noise in factories. The edge DMoA-based PnC can be customized to absorb a desirable amount of ambient energy at the desired frequency leading to improved operating environments.
@@
Note we include nominal viscous damping, $c_{\mathrm{m,r}}=0.01$ and $c_{\mathrm{k,r}}=0$, in numerical simulations presented in this section, to realize bounded acoustic responses.
3.1 PnC with a virtual ground
@@
Figure 6: PnC with a virtual ground. (a-i) Schematics of PnCs featuring a rigid inertial and a non-inertial ground. (a-ii) FFTs of velocity response, $\dot{v}_{\mathrm{d}}(\bar{t})$ in an PnC with rigid inertial grounding (red) and non-inertial grounding (blue). Both PnCs are subject to force $\mathrm{F}(\bar{t})$ (green), however, the rigid grounding case has no environmental noise while the non-inertial grounding case features a broadband environmental noise (orange) contaminating the PnC response via the grounding springs. (b-i) Schematic of a modified PnC where the last mass functions as a virtual ground, and isolates the PnC from undesirable external noise while simultaneously retaining the edge defect mode characteristics relatively intact. (b-ii) FFTs of $\dot{v}_{\mathrm{d}}(\bar{t})$ in an PnC with virtual grounding subject to force $\mathrm{F}(\bar{t})$. (c-i) Reduced order mass-spring model of the PnC with virtual grounding, where grounding springs connect all masses to the last mass (opposite boundary), (c-ii) eigenmode of the PnC at $\Omega_{\mathrm{d,VG}}=0.8447$, and (c-iii) the convergence plot of the edge defect frequency, $\Omega_{\mathrm{d,VG}}$.
@@
As described in Sec. 2, although grounding elements are not necessary for the existence of defect modes, their presence enables the existence of a defect mode as the principal eigenmode of a DMoA-based PnC within the bounded BG – $\Omega\in[0,\sqrt{k_{\mathrm{g,r}}})$. However, PnCs and metamaterials have been deployed in a variety of environments, e.g., rotating shafts[52], aerial vehicles[41], seismic structures[53] where a true inertial frame of reference may not be accessible. In this context, we present a modified edge DMoA-based PnC that can function unhindered without a rigid inertial reference by utilizing one of its boundaries as a virtual ground. Though elastic cross-connections beyond nearest neighbors have been explored in periodic structures[54, 55] exhibiting rotons and statons, the edge defect-embedded PnC with a virtual ground presents a unique mechanically realizable design (see Fig. A4) with simple inter-unit-cell connections, that has no infinitely-periodic counterpart to model the wave dispersion owing to the presence of material defects at a finite boundary.
@@
The schematics in Fig. 6(a,b-i) contrast PnC configurations with an inertial, non-inertial, and virtual ground. Fig. 6(a,b-ii) contrasts the FFTs of the edge defect velocities when subject to a Gaussian wave-packet type forcing, $\mathrm{F}(\bar{t})$ (centered at $\Omega_{\mathrm{d,DMoA}}$) in these three grounding configurations. The FFT (red) in the rigid grounding scenario reveals a behavior consistent with dynamics characterized by the edge DMoA PnC model in Fig. 1b. However, the presence of broadband environmental noise (orange) renders the grounding mechanism non-inertial, causing the FFT (blue) in this scenario to deviate from its rigid grounding counterpart. Even though the environmental noise is chosen to have a maximum amplitude an order lower ($O(10^{-1})$) than the excitation signal, $\mathrm{F}(\bar{t})$, the impact on the PnC dynamics is evident.
@@
Since the nature of the external grounding reference is dictated by the operating environment, we consider an alternate PnC configuration as shown in Fig. 6(b-i) to mitigate the impact of a non-inertial grounding reference. In the modified PnC, all constituent masses are now connected to the last mass which functions as an in-house virtual ground, isolating all PnC masses from the external environmental noise. Fig. 6(c-i) presents an equivalent mass-spring reduced order model for this altered PnC configuration with a virtual ground. An eigen-analysis (see Sec. A6) reveals completely new eigenvalues compared to the edge DMoA-based PnC, with three BG resonances – $\Omega=0$ and $\Omega_{\mathrm{d,VG}}=0.8447$ in the low-frequency BG, and $\Omega=6.4095$ in the high-frequency BG, and the remainder lying within the PB range of the base MoA PnC. Resonances at $\Omega=0$ and $\Omega_{\mathrm{d,VG}}=0.8447$, respectively, correspond to the rigid body translation mode and the localized mode shape plotted in Fig. 6(c-ii) while the resonance at $\Omega=6.4095$ corresponds to a new eigenmode localized at the last mass, created as a result of the altered grounding spring connections. Though the eigenmode frequencies and shapes are significantly different from those of an edge DMoA-based PnC, the eigenmode at $\Omega_{\mathrm{d,VG}}=0.8447$ (Fig. 6(c-ii)) emerges as an exception, existing relatively unaltered compared to the edge defect mode shown in Fig. 1(b-ii), owing to the last mass being a nodal point in the defect mode. The FFT curve (red) in Fig. 6(b-ii) captures the PnC dynamics in the presence of an in-house virtual ground which is identical to that of an PnC with a rigid ground (Fig. 6(a-ii)). Therefore, our modified analytical approach in conjunction with the presence of an in-house virtual ground provides an enhanced capability to design a DMoA-based PnC that can function efficiently at a prescribed BG resonance regardless of the nature of the operating environment.
@@
A closer look at Fig. 6(b-ii) reveals a slight change in defect modal frequency, $\Omega_{\mathrm{d,VG}}$ compared to the analytical DMoA PnC edge defect frequency, $\Omega_{\mathrm{d,DMoA}}=0.8165$. This discrepancy stems from the limitation that the decaying eigenmode reaches a near-zero and not an absolute zero displacement at the last mass, i.e., $v_{\mathrm{N}}\approx 0$. Fig. 6(c-iii) illustrates this phenomenon, where choosing a larger PnC (large $\mathrm{N}$) aids in converging the displacement, $v_{\mathrm{N}}\rightarrow 0$ and the frequency, $\Omega_{\mathrm{d,VG}}\rightarrow\Omega_{\mathrm{d,DMoA}}$.
A slight modification to the 3D structure described in Fig. 1(b-ii) can enable a physical realization of a PnC with a virtual ground. If the grounding frame is rigidly connected to the last aluminum plate (bottom boundary) of the 3D structure, instead of an external grounding reference (see Fig. A4), the frame can essentially be viewed as an extension of the last aluminum plate allowing the grounding clips ($k_{\mathrm{g}}$), to elastically link the remaining plate masses to the virtual ground (last aluminum plate), as depicted in Fig. 6(b-ii).
3.2 Customized acoustic absorber